On the spectrum of second order elliptic differential operators with periodic coefficients

Leonid Friedlander University of Arizona

Under an additional symmetry assumption, we prove that the spectrum of an elliptic second order differential operator with epriodic coefficients in R^{n} is purely absolutely continuous.